Find the exact arc length of the curve


x = cos 3t, y = sin 3t ;(0 ≤ t ≤π)

1
Expert's answer
2021-12-05T17:45:59-0500

"L=\int \sqrt{(x')^2+(y')^2}dt"


"x'=-3sin3t, y'=3cos3t"


"L=\int^{\pi}_0 \sqrt{(-3sin3t)^2+(3cos3t)^2}dt=3\int^{\pi}_0 dt=3\pi"